# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ 4.2. Identifying Dehn surgery coefficients
    Let $M$ be a compact, orientable 3-manifold such that $\partial M$ is a torus. Then there exists a unique slope $c\in\text{slp}(\partial M)$ so that $c$ represents the null homology class in $H_{1}(M;\mathbb{Q})$; equivalently, (any representative of) $c$ belongs to the kernel of the homomorphism $\iota_{\ast}:H_{1}(\partial M;\mathbb{Z})\to H_{1}(M;\mathbb{Z})_{\mathrm{free}}$ induced by inclusion. This boundary slope is called the “rational null-homologous boundary slope” of $M$.
@@
(1) follows from a general argument using Poincaré-Lefschetz duality and the long exact sequence of homology groups, see [34, Lemma 8.15].
For (2), $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ implies that $H_{1}(M_{c};\mathbb{Z})\cong H_{1}(N_{\overline{\psi}(c)};\mathbb{Z})$, so $b_{1}(M_{c})=b_{1}(N_{\overline{\psi}(c)})$. Note that $b_{1}(M_{c_{0}})=b_{1}(M)$, and $b_{1}(M_{c})=b_{1}(M)-1$ if $c\neq c_{0}$. Likewise, $b_{1}(N_{c_{0}^{\prime}})=b_{1}(N)$, and $b_{1}(N_{c^{\prime}})=b_{1}(N)-1$ if $c^{\prime}\neq c_{0}^{\prime}$. Comparing the first Betti number, we obtain that $b_{1}(M)=b_{1}(N)$ and $\overline{\psi}(c_{0})=c_{0}^{\prime}$. ∎
@@
Let $K$ be a knot in $S^{3}$, and let $X_{K}=S^{3}\setminus\mathring{N}(K)$ denote its complement. Dehn fillings of $X_{K}$ can be parametrized by ${}\mathbb{Q}\cup\{\infty\}$, once we fix an orientation of $S^{3}$. To be explicit, let $[m]$ be a slope on $\partial N(K)=\partial X_{K}$, so that the simple closed curve $[m]$ bounds a disk in $N(K)$, which is called the meridian of $K$. By Alexander duality, $\partial X_{K}$ contains a unique null-homologous slope, which we denote as $[l]$, or the preferred longitude of $K$. We further choose orientations on $m$ and $l$, so that the ordered basis $(m,l)$ yields the boundary orientation of $\partial N(K)$, where $N(K)$ inherits its orientation from $S^{3}$ as a submanifold. Then, the boundary slopes of $X_{K}$ can be expressed by $[pm+ql]$, where $(p,q)$ is a pair of coprime integers. The Dehn filling of $X_{K}$ along the slope $[pm+ql]$ is also called a $\frac{p}{q}$-Dehn surgery of $K$, where $\frac{p}{q}\in\mathbb{Q}\cup\{\infty\}$, and is denoted by $K(\frac{p}{q})$. The well-definedness follows from the fact that $[pm+ql]$ and $[-pm-ql]$ represents the same boundary slope.
@@
Let $K$ and $K^{\prime}$ be hyperbolic knots or hyperbolic-type satellite knots in $S^{3}$. Suppose that $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$. Then there exists $\sigma\in\{1,-1\}$, such that for any $r\in\mathbb{Q}\cup\{\infty\}$, $\widehat{\pi_{1}K({r})}\cong\widehat{\pi_{1}K^{\prime}({\sigma r})}$.
@@
For brevity of notation, we denote $M=X_{K}$ and $M^{\prime}=X_{K^{\prime}}$. According to Proposition 2.2, $K$ and $K^{\prime}$ are either both hyperbolic knots, or both satellite knots. By Theorem A for hyperbolic knots and Proposition 3.3 for hyperbolic-type satellite knots, there exists an isomorphism $\psi:\pi_{1}\partial M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial M^{\prime}$ such that for any $c\in\text{slp}(\partial M)$,
@@
Now we show that $\psi(m)=\pm m^{\prime}$. Indeed, $M_{[m]}\cong S^{3}$, and $\widehat{\pi_{1}M^{\prime}_{\overline{\psi}[m]}}\cong\widehat{\pi_{1}M_{[m]}}$ is the trivial group. The same argument as in the proof of Theorem 4.1, i.e. the residual finiteness of the fundamental group together with the validity of Poincaré conjecture, implies that $M^{\prime}_{\overline{\psi}[m]}$ is homeomorphic to $S^{3}$. Note that $K^{\prime}$ is a non-trivial knot; according to [23, Theorem 2], the only Dehn-surgery of $K^{\prime}$ yielding $S^{3}$ is the trivial one, i.e. the $\infty$-surgery. Therefore, $\overline{\psi}[m]$ represents the same boundary slope as $[m^{\prime}]$, i.e. $\psi(m)=\pm m^{\prime}$.
@@
For any coprime pair of integers $(p,q)$, $K(\frac{p}{q})=M_{[pm+ql]}$, and $K^{\prime}(\sigma\frac{p}{q})=M^{\prime}_{[\sigma pm^{\prime}+ql^{\prime}]}=M^{\prime}_{[\delta(\sigma pm^{\prime}+ql^{\prime})]}=M^{\prime}_{\overline{\psi}[pm+ql]}$. Thus, $\widehat{\pi_{1}K(\frac{p}{q})}\cong\widehat{\pi_{1}K^{\prime}(\sigma\frac{p}{q})}$ by (4.6), which finishes the proof of this proposition. ∎
@@ 4.3. Criteria for profinite rigidity of a knot complement
Let $K$ be a knot in $S^{3}$, and $\alpha\in\mathbb{Q}$ be a slope on $\partial N(K)$.
@@
Generally, for $\alpha\neq 0$, neither one of unoriented-characterising and characterising directly implies the other one. However, for the 00-slope, we have the following relation.
@@
Let $K$ be a knot in $S^{3}$. If 00 is a characterising slope of $K$, then 00 is also an unoriented characterising slope of $K$.
@@
We now prove a generalized version of Corollary D. In fact, we can generalize Definition 4.7 (1) to a set of slopes.
@@
Let $N$ be a compact, orientable 3-manifold such that $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}N}$. According to Theorem 4.1 for the hyperbolic case and Theorem 4.3 for the satellite case, $N\cong X_{J}$ is also the complement of a hyperbolic knot or a hyperbolic-type satellite knot $J$. In addition, by Theorem 4.5, there exists $\sigma\in\{\pm 1\}$ such that $\widehat{\pi_{1}K(r)}\cong\widehat{\pi_{1}J(\sigma r)}$ for all $r\in\mathbb{Q}\cup\{\infty\}$. By possibly replacing $J$ with its mirror image, we may assume $\sigma=1$.
Therefore, $\widehat{\pi_{1}K(\alpha)}\cong\widehat{\pi_{1}J(\alpha)}$ for all $\alpha\in\Lambda$. Then, condition (2) implies that $K(\alpha)\cong J(\alpha)$ for all $\alpha\in\Lambda$, and condition (1) thus implies that $J$ is either equivalent or mirror image to $K$. Hence, $N\cong X_{J}\cong X_{K}$. ∎
@@
Now we can apply Corollary D as a technique for the proof of Theorem E.
Theorem E.
@@
    the Pretzel knots $P(-3,3,2n+1)$, where $n\in\mathbb{Z}$ (Figure 7(c));
@@
The knots in (1), (2) and (3) are hyperbolic knots, while the knots in (4) are hyperbolic-type satellite knots. In fact, the $5_{2}$ knot is the $k3_{2}$ hyperbolic knot as tabulated in the census [13], and the $15n_{43522}$ knot is the $k8_{128}$ hyperbolic knot in the census [14]. The hyperbolicity of $P(-3,3,2n+1)$ follows from [36]. The satellite knots $\mathcal{W}^{+}(T_{2,3},2)$ and $\mathcal{W}^{-}(T_{2,3},2)$ are Whitehead-doubles of non-trivial knots, and so they are hyperbolic-type satellite knots.
@@
Therefore, we can adopt Corollary D to prove the profinite rigidity of these knot complements. Indeed, [5, Theorem 1.1] shows that 00 is a characterising slope for $5_{2}$, and [4, Theorem 1.1] shows that 00 is a characterising slope for $15n_{43522}$, $P(-3,3,2n+1)$ or $\mathcal{W}^{\pm}(T_{2,3},2)$. Thus, by Lemma 4.8, 00 is an unoriented characterising slope for any listed knot $K$. Hence, according to Corollary D, it suffices to show that $K(0)$ is profinitely rigid among all closed, orientable 3-manifolds.
@@
Indeed, when $K$ is the hyperbolic knot listed in (1), (2) and (3), Cheetham-West showed in the proof of [16, Theorem 1.1] that $K(0)$ is a graph manifold with JSJ-graph a single vertex and one cycle edge. Therefore, $K(0)$ is profinitely rigid in closed, orientable 3-manifolds according to the profinite classification of closed graph manifolds by Wilkes [49, Theorem A].
@@
We claim that when $K=\mathcal{W}^{\pm}(T_{2,3},2)$, $K(0)$ is a graph manifold with JSJ-graph shown in Figure 9(a). Let $J$ denote the companion knot of $K$, which is a right-handed trefoil. The complement $X_{J}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$. Let $M_{1}$ denote the complement of the Whitehead link, then $X_{K}=M_{1}\cup_{T^{2}}X_{J}$. Note that each boundary component of $M_{1}$ already contains a null-homologous slope, corresponding to the preferred longitude of the Whitehead link. Thus, the 00-surgery on $K$ is in fact the Dehn filling corresponding to the longitude of the pattern Whitehead link. Let $\mathcal{W}(0)$ denote 00-surgery on one component of the Whitehead link. Then, $K(0)=\mathcal{W}(0)\cup_{T^{2}}X_{J}$. In fact, according to [37], $\mathcal{W}(0)$ is a graph manifold obtained from a single Seifert fibered piece $P\times S^{1}$ by gluing together two of its boundary components, where $P=S^{2}\setminus\{3\text{ disks}\}$ denotes a pair of pants. In addition, on $\partial\mathcal{W}(0)$, the slope of a regular fiber in the Seifert fiberation is exactly the meridian of the unfilled Whitehead link component. Since $K=\mathcal{W}^{\pm}(T_{2,3},2)$ is a Whitehead double of twisting number $2$, the meridian of the pattern Whitehead link corresponds to the slope $2$ on $\partial N(J)$. On the other hand, $J=T_{2,3}$, so the slope of a regular fiber of the Seifert fibered space $X_{J}$ on $\partial N(J)$ is $6$ in the ${}\mathbb{Q}\cup\{\infty\}$-parametrization. Therefore, the gluing between $\mathcal{W}(0)$ and $X_{J}$ along $T^{2}=\partial N(J)$ is not fiber-parallel. Thus, $\partial N(J)$ is a JSJ-torus of $X_{K}$ according to [3, Proposition 1.9], and the JSJ-graph of $X_{K}$ is obtained by connecting the JSJ-graphs of $\mathcal{W}(0)$ and $X_{J}$ along a single edge corresponding to $\partial N(J)$, which is exactly shown as Figure 9(a).
@@ 5. Profinite detection of exceptional Dehn fillings
In this section, we focus on cusped hyperbolic 3-manifolds. Let $M$ be a finite-volume hyperbolic 3-manifold with cusps. It is known that most Dehn fillings of $M$ are also hyperbolic [45, Theorem 5.8.2]. A Dehn filling $M_{(c_{i})}$ is called exceptional if $M_{(c_{i})}$ is not a finite-volume hyperbolic 3-manifold.
@@
as the collection of all exceptional slopes on $\partial_{i_{0}}M$. Indeed, [45, Theorem 5.8.2] implies that only finitely many slopes on $\partial_{i_{0}}M$ yield exceptional Dehn fillings, i.e. #$\#\mathcal{E}(M,\partial_{i_{0}}M)<+\infty$.
Note that any finite-volume hyperbolic 3-manifold is aspherical and atoroidal, and thus it does not contain an embedded Klein bottle. Therefore, we consider three particular subsets of $\mathcal{E}(M,\partial_{i_{0}}M)$:
$\begin{gathered}\mathcal{E}_{s}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\mid M_{c}\text{ is not aspherical}\right\};\\ \mathcal{E}_{t}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\mid M_{c}\text{ is toroidal}\right\};\\ \mathcal{E}_{k}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\left|\,\begin{gathered}M_{c}\text{ is aspherical, and }M_{c}\\ \text{ contains an embedded Klein bottle}\end{gathered}\right.\right\}.\end{gathered}$

### v3

@@ 4.2. Identifying Dehn surgery coefficients
    Let $M$ be a compact, orientable 3-manifold such that $\partial M$ is a torus. Then, there exists a unique slope $c\in\text{slp}(\partial M)$ so that $c$ represents the null homology class in $H_{1}(M;\mathbb{Q})$; equivalently, $c$ belongs to the kernel of the homomorphism $\iota_{\ast}:H_{1}(\partial M;\mathbb{Z})\to H_{1}(M;\mathbb{Z})_{\mathrm{free}}$ induced by inclusion. This boundary slope is called the “rational null-homologous boundary slope” of $M$.
@@
(1) follows from the so-called “half lives half dies” theorem, see [34, Lemma 8.15].
For (2), by Lemma 2.1, $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ implies that $H_{1}(M_{c};\mathbb{Z})\cong H_{1}(N_{\overline{\psi}(c)};\mathbb{Z})$, so $b_{1}(M_{c})=b_{1}(N_{\overline{\psi}(c)})$. Note that $b_{1}(M_{c_{0}})=b_{1}(M)$, and $b_{1}(M_{c})=b_{1}(M)-1$ if $c\neq c_{0}$. Likewise, $b_{1}(N_{c_{0}^{\prime}})=b_{1}(N)$, and $b_{1}(N_{c^{\prime}})=b_{1}(N)-1$ if $c^{\prime}\neq c_{0}^{\prime}$. Comparing the first Betti numbers, we derive that $b_{1}(M)=b_{1}(N)$ and $\overline{\psi}(c_{0})=c_{0}^{\prime}$. ∎
@@
Let $K$ be a knot in $S^{3}$, and let $X_{K}=S^{3}\setminus\mathring{N}(K)$ denote its complement. Dehn fillings of $X_{K}$ can be parametrized by $\mathbb{Q}\cup\{\infty\}$, once we fix an orientation of $S^{3}$. To be explicit, let $[m]$ be a slope on $\partial N(K)=\partial X_{K}$, so that the simple closed curve $[m]$ bounds a disk in $N(K)$, which is called the meridian of $K$. By Alexander duality, $\partial X_{K}$ contains a unique null-homologous slope, which we denote as $[l]$, or the preferred longitude of $K$. We further choose orientations on $m$ and $l$, so that the ordered basis $(m,l)$ yields the boundary orientation of $\partial N(K)$, where $N(K)$ inherits its orientation from $S^{3}$ as a submanifold. Then, the boundary slopes of $X_{K}$ can be expressed by $[pm+ql]$, where $(p,q)$ is a pair of coprime integers. The Dehn filling of $X_{K}$ along the slope $[pm+ql]$ is also called a $\frac{p}{q}$-Dehn surgery of $K$, where $\frac{p}{q}\in\mathbb{Q}\cup\{\infty\}$, and is denoted by $K(\frac{p}{q})$. The well-definedness follows from the fact that $[pm+ql]$ and [pmql][-pm-ql] represent the same boundary slope.
@@
Let $K$ and $K^{\prime}$ be hyperbolic knots or hyperbolic-type satellite knots in $S^{3}$. Suppose that $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$. Then, there exists $\sigma\in\{1,-1\}$, such that for any $r\in\mathbb{Q}\cup\{\infty\}$, $\widehat{\pi_{1}K({r})}\cong\widehat{\pi_{1}K^{\prime}({\sigma r})}$.
@@
For brevity of notation, we denote $M=X_{K}$ and $M^{\prime}=X_{K^{\prime}}$. According to Proposition 2.3, $K$ and $K^{\prime}$ are either both hyperbolic knots, or both hyperbolic-type satellite knots. By Theorem A for hyperbolic knots and Theorem 3.7 for hyperbolic-type satellite knots, there exists an isomorphism $\psi:\pi_{1}\partial M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial M^{\prime}$ such that for any $c\in\text{slp}(\partial M)$,
@@
Now we show that $\psi(m)=\pm m^{\prime}$. Indeed, $M_{[m]}\cong S^{3}$, and $\widehat{\pi_{1}M^{\prime}_{\overline{\psi}[m]}}\cong\widehat{\pi_{1}M_{[m]}}$ is the trivial group. The same argument as in the proof of Theorem 4.1, i.e. the residual finiteness of the fundamental group together with the validity of Poincaré conjecture, implies that $M^{\prime}_{\overline{\psi}[m]}$ is homeomorphic to $S^{3}$. Note that $K^{\prime}$ is a non-trivial knot; according to [24, Theorem 2], the only Dehn-surgery of $K^{\prime}$ yielding $S^{3}$ is the trivial one, i.e. the $\infty$-surgery. Therefore, $\overline{\psi}[m]$ represents the same boundary slope as $m^{\prime}]$, i.e. $\psi(m)=\pm m^{\prime}$.
@@
For any coprime pair of integers $(p,q)$, $M_{[pm+ql]}=K(\frac{p}{q})$, and $M^{\prime}_{\overline{\psi}[pm+ql]}=M^{\prime}_{[\delta(\sigma pm^{\prime}+ql^{\prime})]}=M^{\prime}_{[\sigma pm^{\prime}+ql^{\prime}]}=K^{\prime}(\sigma\frac{p}{q})$. Thus, $\widehat{\pi_{1}K(\frac{p}{q})}\cong\widehat{\pi_{1}K^{\prime}(\sigma\frac{p}{q})}$ by (4.6), which finishes the proof of this proposition. ∎
@@ 4.3. Criteria for profinite rigidity of a knot complement
Let $K$ be a knot in $S^{3}$, and let $\alpha\in\mathbb{Q}$ be a slope on $\partial N(K)$.
@@
Generally, for $\alpha\neq 0$, neither one of unoriented-characterising and characterising directly implies the other one. However, for the 0 slope, we have the following relation.
@@
Let $K$ be a knot in $S^{3}$. If 0 is a characterising slope of $K$, then 0 is also an unoriented characterising slope of $K$.
@@
Definition 4.7 (1) can also be generalized to a set of slopes.
@@
Let $N$ be a compact, orientable 3-manifold such that $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}N}$. According to Theorem 4.1 for the hyperbolic case and Theorem 4.3 for the satellite case, $N\cong X_{J}$ is also the complement of a hyperbolic knot or a hyperbolic-type satellite knot $J$. In addition, by Theorem 4.5, there exists $\sigma\in\{\pm 1\}$ such that $\widehat{\pi_{1}K(r)}\cong\widehat{\pi_{1}J(\sigma r)}$ for all $r\in\mathbb{Q}\cup\{\infty\}$. By possibly replacing $J$ with its mirror image, we may assume $\sigma=1$, so $\widehat{\pi_{1}K(\alpha)}\cong\widehat{\pi_{1}J(\alpha)}$ for all $\alpha\in\Lambda$. Then, condition (2) implies that $K(\alpha)\cong J(\alpha)$ for all $\alpha\in\Lambda$, and condition (1) thus implies that $J$ is either equivalent or mirror image to $K$. Hence, $N\cong X_{J}\cong X_{K}$. ∎
@@
Now we can apply Corollary D as a technique for the proof of Corollary E.
Corollary E.
@@
    the Pretzel knots P(3,3,2n+1)P(-3,3,2n+1), where $n\in\mathbb{Z}$ (Figure 7(c));
@@
The knots in (1), (2) and (3) are hyperbolic knots, while the knots in (4) are hyperbolic-type satellite knots. In fact, the $5_{2}$ knot is the $k3_{2}$ hyperbolic knot as tabulated in the census [13], and the $15n_{43522}$ knot is the $k8_{128}$ hyperbolic knot in the census [15]. The hyperbolicity of P(3,3,2n+1)P(-3,3,2n+1) follows from [36]. The satellite knots $\mathcal{W}^{+}(T_{2,3},2)$ and $\mathcal{W}^{-}(T_{2,3},2)$ are Whitehead-doubles of non-trivial knots, and so they are hyperbolic-type satellite knots.
@@
Therefore, we can apply Corollary D to prove the profinite rigidity of these knot complements. Indeed, [5, Theorem 1.1] shows that 0 is a characterising slope for $5_{2}$, and [4, Theorem 1.1] shows that 0 is a characterising slope for $15n_{43522}$, P(3,3,2n+1)P(-3,3,2n+1) or $\mathcal{W}^{\pm}(T_{2,3},2)$. Thus, by Lemma 4.8, 0 is an unoriented characterising slope for any listed knot $K$. Hence, according to Corollary D, it suffices to show that $K(0)$ is profinitely rigid among all closed, orientable 3-manifolds.
@@
Indeed, when $K$ is one of the hyperbolic knots listed in (1), (2) and (3), Cheetham-West showed in the proof of [17, Theorem 1.1] that $K(0)$ is a graph manifold with JSJ-graph a single vertex and one cycle edge. Therefore, $K(0)$ is profinitely rigid in closed, orientable 3-manifolds according to the profinite classification of closed graph manifolds by Wilkes [49, Theorem A].
@@
We claim that when $K=\mathcal{W}^{\pm}(T_{2,3},2)$, $K(0)$ is a graph manifold with JSJ-graph shown in Figure 9(a). Let $J$ denote the companion knot of $K$, which is a right-handed trefoil. The complement $X_{J}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$. Let $M_{1}$ denote the complement of the Whitehead link, then $X_{K}=M_{1}\cup_{T^{2}}X_{J}$. Note that each boundary component of $M_{1}$ already contains a null-homologous slope, corresponding to the preferred longitude of the Whitehead link. Thus, the 0-surgery on $K$ is in fact the Dehn filling corresponding to the longitude of the pattern Whitehead link. Let $\mathcal{W}(0)$ denote 0-surgery on one component of the Whitehead link. Then, $K(0)=\mathcal{W}(0)\cup_{T^{2}}X_{J}$. In fact, according to [37], $\mathcal{W}(0)$ is a graph manifold obtained from a single Seifert fibered piece $P\times S^{1}$ by gluing together two of its boundary components, where $P=S^{2}\setminus\{3\text{ disks}\}$ denotes a pair of pants. In addition, on $\partial\mathcal{W}(0)$, the slope of a regular fiber in the Seifert fibration is exactly the meridian of the unfilled Whitehead link component. Since $K=\mathcal{W}^{\pm}(T_{2,3},2)$ is a Whitehead double of twisting number $2$, the meridian of the pattern Whitehead link corresponds to the slope $2$ on $\partial N(J)$. On the other hand, $J=T_{2,3}$, so the slope of a regular fiber of the Seifert fibered space $X_{J}$ on $\partial N(J)$ is $6$ in the $\mathbb{Q}\cup\{\infty\}$-parametrization. Therefore, the gluing between $\mathcal{W}(0)$ and $X_{J}$ along $T^{2}=\partial N(J)$ is not fiber-parallel. Thus, $\partial N(J)$ is a JSJ-torus of $X_{K}$ according to [3, Proposition 1.9], and the JSJ-graph of $X_{K}$ is obtained by connecting the JSJ-graphs of $\mathcal{W}(0)$ and $X_{J}$ along a single edge corresponding to $\partial N(J)$, which is exactly shown as Figure 9(a).
@@ 5. Profinite detection of exceptional Dehn fillings
In this section, we focus on general cusped hyperbolic 3-manifolds. Let $M$ be a finite-volume hyperbolic 3-manifold with cusps. By Thurston’s theorem [45, Theorem 5.8.2], most Dehn fillings of $M$ are also hyperbolic. A Dehn filling $M_{(c_{i})}$ is called exceptional if $M_{(c_{i})}$ is not a finite-volume hyperbolic 3-manifold.
@@
as the collection of all exceptional slopes on $\partial_{i_{0}}M$. Indeed, [45, Theorem 5.8.2] implies that only finitely many slopes on $\partial_{i_{0}}M$ yield exceptional Dehn fillings, i.e. $\mathcal{E}(M,\partial_{i_{0}}M)<+\infty$.
Note that any finite-volume hyperbolic 3-manifold is aspherical and atoroidal, and it does not contain an embedded Klein bottle. Therefore, we consider three particular subsets of $\mathcal{E}(M,\partial_{i_{0}}M)$:
[…]